OpenStudy (anonymous):

Which has a greater area, four circles, each with radius 1 m, or one circle with radius 4 m? Explain , Already know the answer the greater one is the one circle but i need help with this question, dont have a compass :( To check your answer, use a compass to draw the large and small circles. Draw a circle with radius 4 cm to represent the large circle. Can you draw the four smaller circles so that they fit inside the large circle, without overlapping?

OpenStudy (johnweldon1993):

Well you DO have the correct answer...just dont do the drawing portion...prove it with the formula for area of a circle

OpenStudy (johnweldon1993):

Area = pi * r^2 so 4 smaller ones = 4( pi * 1^2) = 4pi 1 bigger one = pi * 4^2 = 16pi so clearly you have the correct answer